# Diffusion Models at the Drug Discovery Frontier: A Review on Generating Small Molecules versus Therapeutic Peptides

### v1

@@ 3.3 Property-Based Ligand Design and Optimization
This area focuses on generating molecules that satisfy multiple objectives simultaneously, such as high binding affinity, favorable drug-likeness (quantified by metrics like QED, Quantitative Estimate of Drug-likeness [130]), appropriate lipophilicity (logP), low toxicity, high membrane permeability, and synthetic accessibility (SA). This multi-objective optimization is often achieved by combining diffusion models with property predictors to guide the generation toward desired regions of the chemical space [131, 132, 133, 134, 103, 53, 135, 58, 40].
@@
Several approaches have been developed for property-guided generation. Conditional diffusion models can be trained to generate molecules with specific property values by conditioning on target property vectors [53]. These models have demonstrated the ability to generate molecules with specified molecular weight, logP, and number of hydrogen bond donors/acceptors while maintaining reasonable accuracy in property prediction [53]. Guidance-based methods employ pre-trained property predictors to steer the diffusion sampling process toward molecules with desired properties [131, 132, 40]. By computing gradients of property predictors with respect to the molecular representation, these methods can navigate the chemical space to optimize multiple properties simultaneously [53, 58].
@@
Multi-objective optimization remains challenging due to the often conflicting nature of different objectives. For instance, increasing lipophilicity (logP) to improve membrane permeability may simultaneously decrease aqueous solubility and increase toxicity risk. Recent work has explored Pareto-optimal generation and multi-objective frameworks, where diffusion models are trained to generate diverse molecules along the Pareto front of multiple objectives [133, 135], allowing medicinal chemists to select candidates that best balance different requirements. Reinforcement learning-guided approaches have also been developed to dynamically balance competing objectives during the generation process [133, 134], while dual diffusion architectures enable simultaneous optimization across multiple design criteria [103].
@@
However, ensuring the synthesizability of the generated molecules remains a major and persistent challenge in the field. While diffusion models can generate chemically valid molecules (as determined by valence rules and RDKit sanitization), these molecules may be synthetically inaccessible or require prohibitively complex synthetic routes. Synthetic accessibility scores (SA scores [136]) provide a rough estimate [131, 134], but they do not guarantee that a practical synthesis route exists. Recent efforts have focused on incorporating retrosynthesis models into the generation process, either by using retrosynthesis feasibility as an additional objective [137] or by generating molecules in a retrosynthetically-aware manner, building molecules from commercially available building blocks through known reaction templates [138, 139]. Alternative approaches evaluate synthesizability by combining retrosynthetic planning with forward reaction prediction to verify route feasibility [140]. Methods that optimize molecular geometry and structural stability have also been proposed to improve the practical viability of generated candidates [58]. Despite these advances, the gap between computational generation and experimental synthesis remains a critical bottleneck [108, 141]—a synthesis barrier that has been identified as a major challenge limiting the real-world impact of generative AI in pharmaceutical development [142]. Bridging this gap by integrating generative models with retrosynthesis prediction and automated experimental validation remains a central goal for the field [143], a challenge shared across modalities, where the synthetic accessibility hurdle for small molecules finds its critical counterpart in the biological stability and production challenges inherent to therapeutic peptides (Section 4).
@@ 4.1 Datasets and Benchmarks for Peptide Design
For specific peptide design tasks, specialized datasets exist. The Antimicrobial Peptide Database (APD3) contains approximately 3,000 experimentally validated antimicrobial peptides with activity data (MIC values, target organisms) [153]. The Database of Antimicrobial Activity and Structure of Peptides (DBAASP) contains over 15,000 entries with detailed activity annotations [154]. For cell-penetrating peptides, CPPsite contains approximately 1,800 entries [155, 156]. However, these specialized datasets are much smaller than those available for small molecules, and the activity data is often heterogeneous (measured under different conditions, against different organisms), making it challenging to train robust predictive models.
@@ 4.2 Generation of Functional Peptide Sequences
Discrete diffusion models for peptide sequences operate by gradually corrupting amino acid sequences through a process of random token replacement or masking, then learning to reverse this process [35]. Several noise processes have been proposed, including uniform transition matrices (where each amino acid has equal probability of transitioning to any other) [35, 66], absorbing state models (where tokens are gradually replaced by a special MASK token) [35], and learned transition matrices that respect amino acid similarity (e.g., hydrophobic amino acids are more likely to transition to other hydrophobic amino acids) [35].
@@ 4.3 Structure-Guided De Novo Peptide Design
A more ambitious goal is to directly generate peptides that fold into specific 3D structures or bind to target protein surfaces with high affinity and specificity. This includes not only linear peptides but also larger, structurally-defined mini-proteins that function as peptide mimetics. This task requires modeling both sequence and structure simultaneously, since the sequence must be compatible with the desired fold and the structure must be stable and functional [180, 181]. Recent deep learning advances, particularly diffusion-based methods, have made significant progress toward achieving this goal [182, 183].
@@
RFdiffusion has demonstrated remarkable success in designing mini-protein binders, a breakthrough that directly paves the way for creating structurally-defined peptides with high efficacy [69]. When tasked with designing binders to challenging protein targets such as influenza hemagglutinin, for instance, RFdiffusion generates backbones that, after sequence design using ProteinMPNN [184], achieve experimental binding affinities in the nanomolar range (e.g., a KD of 28 nM for an influenza binder) in approximately 19% of tested designs [69]. This success rate is substantially higher than previous computational design methods, which typically achieved success rates below 5% [185, 186]. The designed binders often exhibit novel folds not present in natural proteins, demonstrating the model’s ability to explore diverse and novel structural topologies within the protein fold space [69]. Furthermore, the approach has been successfully extended to designing high-affinity binders for challenging helical peptide targets, yielding picomolar to sub-nanomolar affinities [187].
@@
Despite these successes, significant challenges unique to peptide therapeutics remain. Generated peptides must be engineered for proteolytic stability to overcome their inherently short in vivo half-lives, a consideration often addressed by incorporating non-canonical amino acids or cyclization, which are not yet fully integrated into diffusion workflows [189]. Furthermore, minimizing potential immunogenicity by avoiding T-cell epitopes is a critical design constraint that requires sophisticated predictive modeling [190]. Ultimately, ensuring that the designed sequence not only folds into the intended conformation but also remains stable and avoids aggregation is paramount, as current models may not fully capture the subtle side-chain interactions governing these properties [191]. Integrating these complex biological and biophysical constraints into the next generation of generative models represents a critical frontier for the field.
@@ 5 Comparison, Challenges, and Future Perspectives
Figure 2: Contrasting Design Paradigms for Small Molecules and Therapeutic Peptides with Diffusion Models. The figure illustrates the distinct challenges and tailored AI-driven solutions for small molecules (left column, a, c, e, g) versus therapeutic peptides (right column, b, d, f, h). (a, b) The primary challenge for small molecules is navigating the vast, discrete chemical space, whereas for peptides, it is conquering the continuous conformational space to achieve a stable fold. (c, d) Consequently, diffusion models are employed for structure-based generation to fit small molecules into binding pockets, while for peptides, they perform structure-guided design by decorating a predefined scaffold. (e, f) Key downstream hurdles also differ: ensuring chemical synthesizability for small molecules versus achieving biological stability against degradation for peptides. (g, h) Finally, solutions are modality-specific: integrating chemical knowledge (e.g., reaction rules) to guide synthesis for small molecules, and engineering stability in peptides through modifications like cyclization or using non-canonical amino acids.
@@
Chemical Space Vast & Discontinuous ($\sim 10^{60}$[11, 15, 192, 12]
@@ 5.2 Shared Hurdles and Common Challenges
Consequently, a critical imperative for the field is to ”close the loop” by integrating generative models with automated experimental validation in a Design-Build-Test-Learn (DBTL) cycle, as illustrated in Figure 3 [222, 223, 224]. Without such a framework, which is now becoming feasible through advances in laboratory automation [225, 226, 227, 228], the design process remains a slow, sequential, and inefficient endeavor [229, 230, 231, 232]. Finally, even with better data and validation, the issue of generalization persists. Like all machine learning models, diffusion models risk overfitting to their training distribution, potentially failing to generate effective and novel solutions for new biological targets or chemical spaces that lie outside their learned domain [233, 131, 234, 235]. Overcoming these interconnected challenges is essential to translate the theoretical promise of diffusion models into tangible therapeutic breakthroughs [236, 237].
@@ 5.3 Future Outlook and Opportunities
The field of diffusion models for drug discovery is rapidly evolving, with future work poised to address current limitations and unlock transformative capabilities. A key frontier is the development of unified frameworks—so-called ”foundation models” for molecular science—that could seamlessly design not only small molecules and peptides but also complex hybrid therapeutics like peptide-drug conjugates (PDCs) and PROTACs from a single, powerful architecture. Enhancing model reliability is also paramount; this involves a shift from ’black box’ generators to interpretable and controllable tools that empower expert-guided design, while integrating first-principles simulations from quantum chemistry and physics to ensure the physical realism of generated candidates. Ultimately, the successful translation of these technologies will hinge on fully realizing the automated Design-Build-Test-Learn (DBTL) paradigm, which promises to accelerate discovery cycles from months to days. This acceleration, however, must be navigated alongside the establishment of clear ethical and regulatory frameworks to guide AI-designed therapeutics safely from concept to clinic.
6 Conclusion
Diffusion models have emerged as a powerful, unified generative framework, demonstrating remarkable versatility in designing both small molecules and therapeutic peptides. While successful in generating novel candidates for both modalities, the path to clinical translation is defined by distinct, fundamental hurdles: for small molecules, the challenge lies in bridging the gap from computational validity to practical chemical synthesizability; for peptides, it is ensuring that de novo structural designs achieve in vivo biological stability and function. The full potential of this technology will be significantly accelerated by closing the Design-Build-Test-Learn loop through deep integration with laboratory automation, which will enable rapid, data-driven iteration. By overcoming these challenges, diffusion models hold the promise to catalyze a fundamental shift in drug discovery—moving from the passive exploration of existing chemical space to the active, purpose-driven creation of novel medicines.
@@ 6 Conclusion
Not applicable.

### v2

@@ 2.3 Property-Based Ligand Design and Optimization
This area focuses on generating molecules that satisfy multiple objectives simultaneously, such as high binding affinity, favorable drug-likeness measured by metrics like the Quantitative Estimate of Drug-likeness (QED) [126], where values closer to 1 suggest a better drug-like profile, appropriate lipophilicity (logP), low toxicity, high membrane permeability, and synthetic accessibility (SA). To achieve this multi-objective optimization, several property-guided generation approaches have been developed. Conditional diffusion models, for example, are trained to generate molecules with specific property values by directly conditioning on target property vectors [48]. These models can produce molecules with specified molecular weight, logP, and hydrogen bond donor/acceptor counts with reasonable accuracy [48]. Alternatively, guidance-based methods employ pre-trained property predictors to steer the diffusion sampling process at inference time [127, 128, 35]. By computing the gradients of property predictors with respect to the molecular representation, these techniques can navigate the chemical space to optimize multiple properties simultaneously [53].
@@
However, optimizing for multiple, often conflicting, objectives remains a significant challenge. For instance, increasing lipophilicity (logP) to improve membrane permeability may concurrently decrease aqueous solubility and heighten toxicity risk. To address this, recent work has explored more sophisticated frameworks. Some studies focus on generating diverse molecules along the Pareto front, providing a set of candidates that represent different trade-offs between objectives [129]. Other advanced strategies include using reinforcement learning to dynamically balance competing goals [130, 131] and developing dual diffusion architectures for simultaneous optimization across multiple design criteria [99].
@@
However, ensuring the synthesizability of the generated molecules remains a major and persistent challenge in the field. While diffusion models can generate chemically valid molecules (as determined by valence rules and RDKit sanitization), these molecules may be synthetically inaccessible or require prohibitively complex synthetic routes. Synthetic accessibility scores (SA scores [132]) provide a rough estimate [127, 131], but they do not guarantee that a practical synthesis route exists. Recent efforts have focused on incorporating models for retrosynthesis, a computational technique for planning chemical synthesis by working backward from the target molecule, into the generation process, either by using retrosynthesis feasibility as an additional objective [133] or by generating molecules in a retrosynthetically aware manner, building molecules from commercially available building blocks through known reaction templates [134, 135]. Alternative approaches evaluate synthesizability by combining retrosynthetic planning with forward reaction prediction to verify route feasibility [136]. Methods that optimize molecular geometry and structural stability have also been proposed to improve the practical viability of generated candidates [53]. Despite these advances, the gap between computational generation and experimental synthesis remains a critical bottleneck [104, 137]—a synthesis barrier that has been identified as a major challenge limiting the real-world impact of generative AI in pharmaceutical development [138]. Bridging this gap by integrating generative models with retrosynthesis prediction and automated experimental validation remains a central goal for the field [139], a challenge shared across modalities, where the synthetic accessibility hurdle for small molecules finds its critical counterpart in the biological stability and production challenges inherent to therapeutic peptides (Section 3).
@@ 3.1 Datasets and Benchmarks for Peptide Design
For specific peptide design tasks, specialized datasets exist. The Antimicrobial Peptide Database (APD3) contains approximately 3000 experimentally validated antimicrobial peptides with activity data (MIC values, target organisms) [149]. The Database of Antimicrobial Activity and Structure of Peptides (DBAASP) contains over 15,000 entries with detailed activity annotations [150]. For cell-penetrating peptides, CPPsite contains approximately 1800 entries [151, 152]. However, these specialized datasets are much smaller than those available for small molecules. Furthermore, they suffer from significant data heterogeneity, a challenge that directly impacts model reproducibility and the creation of reliable benchmarks. For instance, antimicrobial activity measured as Minimum Inhibitory Concentration (MIC) can vary dramatically depending on the bacterial strain and assay protocol, while cell-penetrating efficiency lacks a universally accepted standard metric. This inconsistency makes it difficult to harmonize data for training robust, generalizable predictive models and underscores the critical need for community-wide standards in peptide bioactivity reporting.
@@ 3.2 Generation of Functional Peptide Sequences
Discrete diffusion models for peptide sequences operate by gradually corrupting amino acid sequences through a process of random token replacement or masking, then learning to reverse this process. The foundational work in this area proposed several noise processes, including uniform transition matrices, absorbing state models, and learned transition matrices that respect amino acid similarity [61]. The uniform transition approach, for instance, has been applied in subsequent peptide generation models [62]. The choice among these noise processes carries significant practical implications for peptide design. The uniform transition matrix, while the simplest to implement, disregards the inherent biochemical similarities between amino acids, treating a transition from Alanine to Valine (both hydrophobic) the same as one to Lysine (hydrophilic). The absorbing state model is particularly well-suited for tasks like sequence inpainting or constrained generation, as the MASK token provides a clear demarcation between fixed regions and those to be generated. Finally, learned transition matrices offer the most sophisticated approach, allowing the model to incorporate prior knowledge, such as amino acid substitution matrices (e.g., BLOSUM), which can potentially improve learning efficiency and generate more biologically plausible intermediate sequences.
@@ 3.3 Structure-Guided De Novo Peptide Design
A more ambitious goal is to directly generate peptides that fold into specific 3D structures or bind to target protein surfaces with high affinity and specificity. This includes not only linear peptides but also larger, structurally defined mini-proteins that function as peptide mimetics. This task requires modeling both sequence and structure simultaneously, since the sequence must be compatible with the desired fold and the structure must be stable and functional [176, 177]. Recent deep learning advances, particularly diffusion-based methods, have made significant progress toward achieving this goal [178, 179].
@@
RFdiffusion has demonstrated remarkable success in designing mini-protein binders, a breakthrough that directly paves the way for creating structurally defined peptides with high efficacy [65]. When tasked with designing binders to challenging protein targets such as influenza hemagglutinin, for instance, RFdiffusion generates backbones that, after sequence design using ProteinMPNN [180], achieve experimental binding affinities in the nanomolar range (e.g., a KD of 28 nM for an influenza binder) in approximately 19% of tested designs [65]. This success rate is substantially higher than previous computational design methods, which typically achieved success rates below 5% [181, 182]. The designed binders often exhibit novel folds not present in natural proteins, demonstrating the model’s ability to explore diverse and novel structural topologies within the protein fold space [65]. Furthermore, the approach has been successfully extended to designing high-affinity binders for challenging helical peptide targets, yielding picomolar to sub-nanomolar affinities [183].
@@
Despite these successes, significant challenges unique to peptide therapeutics remain. Generated peptides must be engineered for proteolytic stability to overcome their inherently short in vivo half-lives, a consideration often addressed by incorporating non-canonical amino acids or cyclization, which are not yet fully integrated into diffusion workflows [185]. Furthermore, minimizing potential immunogenicity by avoiding T-cell epitopes—specific peptide fragments recognized by the immune system that can trigger an unwanted response—is a critical design constraint that requires sophisticated predictive modeling [186]. Ultimately, ensuring that the designed sequence not only folds into the intended conformation but also remains stable and avoids aggregation is paramount, as current models may not fully capture the subtle side-chain interactions governing these properties [187]. Integrating these complex biological and biophysical constraints into the next generation of generative models represents a critical frontier for the field.
While the specific bottlenecks differ, the parallel evolution of diffusion models in these two domains invites a systematic comparison. Synthesizing these distinct modality-specific insights is essential to identify the shared fundamental challenges and to envision the future trajectory of the field toward a unified, automated discovery paradigm.
@@ 4 Comparison, Challenges, and Future Perspectives
Chemical Space Vast & Discontinuous ($\sim$$0^{60}[11$, 15, 188, 12]
@@
Figure 2: Contrasting Design Paradigms for Small Molecules and Therapeutic Peptides with Diffusion Models. The figure illustrates the distinct challenges and tailored AI-driven solutions for small molecules (left column, a,c,e,g) versus therapeutic peptides (right column, b,d,f,h). (a,b) The primary challenge for small molecules is navigating the vast, discrete chemical space, whereas for peptides, it is conquering the continuous conformational space to achieve a stable fold. (c,d) Consequently, diffusion models are employed for structure-based generation to fit small molecules into binding pockets, while for peptides, they perform structure-guided design by decorating a predefined scaffold. (e,f) Key downstream hurdles also differ: ensuring chemical synthesizability for small molecules versus achieving biological stability against degradation for peptides. (g,h) Finally, solutions are modality-specific: integrating chemical knowledge (e.g., reaction rules) to guide synthesis for small molecules, and engineering stability in peptides through modifications like cyclization or using non-canonical amino acids. Explanation of symbols: The red crosses (X) indicate synthetic infeasibility (e) or blocked enzymatic degradation (f, h). In (e), the colored spheres represent atoms within a complex molecular graph structure.
@@ 4.2 Shared Hurdles and Common Challenges
Consequently, a critical imperative for the field is to “close the loop” by integrating generative models with automated experimental validation in a Design-Build-Test-Learn (DBTL) cycle, as illustrated in Figure 3 [218, 219, 220]. The implementation of such a cycle creates a direct pathway from in silico hypothesis generation to experimental validation and back, enabling a rapid, iterative flow where data from one round directly informs the next. Without such a framework, which is now becoming feasible through advances in laboratory automation [221, 222, 223, 224], the design process remains a slow, sequential, and inefficient endeavor [225, 226, 227, 228]. Finally, even with better data and validation, the issue of generalization persists. Like all machine learning models, diffusion models risk overfitting to their training distribution, potentially failing to generate effective and novel solutions for new biological targets or chemical spaces that lie outside their learned domain [229, 127, 230, 231]. Overcoming these interconnected challenges is essential to translate the theoretical promise of diffusion models into tangible therapeutic breakthroughs [232, 233].
@@ 4.3 Future Outlook and Opportunities
A critical measure of success for any drug discovery technology is its path to clinical translation. While molecules designed specifically by diffusion models have not yet entered human clinical trials, the broader field of generative AI provides a strong and encouraging precedent. Companies such as Insilico Medicine, Exscientia, and Recursion Pharmaceuticals have successfully advanced AI-designed small molecules into various stages of clinical development, validating the principle that AI can indeed yield viable therapeutic candidates [234, 235]. This industry progress sets the stage for diffusion models, which represent the state-of-the-art in generative capability. Several leading biotechnology companies are now deeply integrating these models into their R&D pipelines, and although much of this work remains proprietary, early reports indicate that diffusion-generated candidates are demonstrating excellent activity and favorable properties in preclinical studies. This marks a rapid transition of the technology from academic exploration to industrial application.
The field of diffusion models for drug discovery is rapidly evolving, with future work poised to address current limitations and unlock transformative capabilities. A key frontier is the development of unified frameworks—so-called “foundation models” for molecular science—that could seamlessly design not only small molecules and peptides but also complex hybrid therapeutics like peptide-drug conjugates (PDCs) and PROTACs from a single, powerful architecture. Enhancing model reliability is also paramount; this involves a shift from ‘black box’ generators to interpretable and controllable tools that empower expert-guided design, while integrating first-principles simulations from quantum chemistry and physics to ensure the physical realism of generated candidates. Ultimately, the successful translation of these technologies will hinge on fully realizing the automated Design-Build-Test-Learn (DBTL) paradigm, as illustrated in Figure 3, which promises to accelerate discovery cycles from months to days. This acceleration, however, must be navigated alongside the establishment of clear ethical and regulatory frameworks to guide AI-designed therapeutics safely from concept to clinic.
5 Conclusions
Diffusion models have emerged as a powerful, unified generative framework, demonstrating remarkable versatility in designing both small molecules and therapeutic peptides. While successful in generating novel candidates for both modalities, the path to clinical translation is defined by distinct, fundamental hurdles: for small molecules, the challenge lies in bridging the gap from computational validity to practical chemical synthesizability; for peptides, it is ensuring that de novo structural designs achieve in vivo biological stability and function. Crucially, the progress of AI-designed drugs now entering clinical trials provides a strong tailwind for the field, validating the potential of these advanced generative approaches. The full potential of this technology will be significantly accelerated by closing the Design-Build-Test-Learn loop through deep integration with laboratory automation, which will enable rapid, data-driven iteration. By overcoming these challenges, diffusion models hold the promise to catalyze a fundamental shift in drug discovery—moving from the passive exploration of existing chemical space to the active, purpose-driven creation of novel medicines.
@@ 5 Conclusions
This work was supported by the Natural Science Foundation of Xinjiang Uygur Autonomous Region (Grant Number: 2024D01C216) and the “Tianchi Talents” introduction plan.